# Self-Speculative Masked Diffusions

### v1

@@ 3.3 Sampling Procedure
We sample our hybrid architecture using Algorithm 2. The general procedure follows Algorithm 1, revealing tokens in update steps until the sequence is complete. In a given update step (with $i$ tokens revealed), the non-causal blocks first generate draft tokens for all unknown positions, $\hat{\bm{x}}^{\sigma(i+1:D)}\sim\overset{\leftrightarrow}{p}_{\theta}(\hat{\bm{x}}^{\sigma(i+1:D)}|\theta(\bm{x}^{\sigma(1:i)}))$. The causal blocks then give target probabilities for the drafted tokens $\{\overset{\rightarrow}{p}_{\theta,\phi}(\hat{\bm{x}}^{\sigma(d)}|\theta(\bm{x}^{\sigma(1:i)}),\phi(\hat{\bm{x}}^{\sigma(i+1:d-1)})\}_{d=i+1}^{D}$. Finally, a speculative sampling inner loop accepts some of the drafted tokens. Appendix B details a more general windowed procedure.
@@
Our procedure bears some similarity to self-speculative (Zhang et al.,, 2024; Elhoushi et al.,, 2024; Liu et al.,, 2024) and Medusa (Cai et al.,, 2024) style approaches where an initial subset of the network is used to parameterize a draft distribution which is then verified by the full transformer. However, using self-speculation for MDMs causes additional complexity because the target distribution given by Equation 6 changes depending on the number of revealed tokens. Consider the targets $\overset{\rightarrow}{p}_{\theta,\phi}(\bm{x}^{\sigma(i+1:D)}|\theta(\bm{x}^{\sigma(1:i)}),\phi)$ and $\overset{\rightarrow}{p}_{\theta,\phi}(\bm{x}^{\sigma(j+1:D)}|\theta(\bm{x}^{\sigma(1:j)}),\phi)$ for $j>i$. The factors for a fixed $d$ differ; i.e., importantly, we have $\overset{\rightarrow}{p}_{\theta,\phi}(\bm{x}^{\sigma(d)}|\theta(\bm{x}^{\sigma(1:i)}),\phi(\bm{x}^{\sigma(i+1:d-1)}))\neq\overset{\rightarrow}{p}_{\theta,\phi}(\bm{x}^{\sigma(d)}|\theta(\bm{x}^{\sigma(1:j)}),\phi(\bm{x}^{\sigma(j+1:d-1)}))$. This is because the non-causal blocks operate on different inputs (one sequence with $i$ revealed tokens, the other with $j$). So the hidden states passed to the causal blocks differ, causing the target distribution for position $\sigma(d)$ to change throughout the generation trajectory. This is not an issue for existing self-speculative models restricted to a single left-to-right ordering. All transformer blocks are causal in this case and never observe future tokens, meaning the target remains unchanged during generation.
@@ 3.4 Theoretical Results
We want to understand here the distribution of samples generated by Algorithm 2 for a given ordering, $p_{\theta,\phi}(\bm{x}^{\sigma(1:D)}|\sigma)$. The target of the inner speculative sampling loop, Equation 6, depends on the current number of revealed tokens, $i$. This therefore implies the need for a combinatorial sum over all possible numbers of tokens revealed during each inner loop. Instead, we show in Proposition 3.1 that $p_{\theta,\phi}(\bm{x}^{\sigma(1:D)}|\sigma)$ can be tractably calculated with $D$ forward passes and $O(D^{2})$ operations using a recursive decomposition. The proof is provided in Appendix C.1.
@@
Consider the sampling scheme defined by Algorithm 2. For a given ordering $\sigma$, let $A^{\sigma(d)}$ denote the event that the token in position $\sigma(d)$ was accepted and let $R^{\sigma(d)}$ denote the event it was rejected and resampled. The distribution of samples output by Algorithm 2, $p_{\theta,\phi}(\bm{x}^{\sigma(1:D)}|\sigma)$, then admits the following decomposition
@@
We can now obtain an Evidence Lower BOund (ELBO) on $\log p_{\theta,\phi}(\bm{x})$,
@@ 3.5 Sampling Improvements
Algorithm 2 assumes each sampling inner loop corresponds to one forward pass of the non-causal blocks and a single verification pass through the remaining causal blocks. We can vary the sample quality-efficiency trade-off by running the speculative inner loop multiple times per single forward pass of the non-causal blocks. The non-causal forward pass first creates a draft $\overset{\leftrightarrow}{p}_{\theta}(\bm{x}^{\sigma(i+1:D)}|\theta(\bm{x}^{\sigma(1:i)}))$ distribution. Draft tokens are then sampled and we compute the target probabilities on these draft tokens using the causal layers. One speculative sampling loop is run to accept a number of these draft tokens finishing with a rejection and a resampling of a token. In the next speculative sampling inner loop we re-use the non-causal hidden states but recompute the target probabilities on the remaining draft tokens, noting the resampling operation will change the target probabilities for subsequent tokens compared to the first speculative sampling inner loop. This procedure is repeated for a user specified number of inner loops, the full algorithm is described in Appendix B. In our experiments, the vast majority of the network is non-causal, so this procedure greatly increases efficiency. To know when to stop generating tokens with the same draft $\overset{\leftrightarrow}{p}_{\theta}(\bm{x}^{\sigma(i+1:D)}|\theta(\bm{x}^{\sigma(1:i)}))$, we use windows (Leviathan et al.,, 2023; Chen et al.,, 2023) to set a maximum number of tokens we can reveal. We discuss window choices in Appendix D.
@@ 4 Related Work
Several speculative sampling works have explored using a single network for both the draft model and target model (Zhang et al.,, 2024; Elhoushi et al.,, 2024; Liu et al.,, 2024), termed self-speculative decoding. These differ from our work by using causal draft and target layers in a strict left-to-right setting. In the same setting, non-causal factorized predictions were explored by Cai et al., (2024) using a mostly causal stack with single-layer factorized prediction heads to draft tokens. We flip this: our network is mostly non-causal with a small final causal layer. Finally, a non-causal draft model and entirely separate causal target model was explored for audio generation in Ziv et al., (2024). We combine these two into a single architecture reducing deployment complexity.
Causal models for any-order AR training were explored by Yang et al., (2019), who introduced XLNet, a specialized two-stream architecture enforcing correct causal dependencies for different orderings. Further architectures used a standard single stream causal architecture and input ordering information through double positional encodings (Pannatier et al.,, 2024) or extra position tokens (Pang et al.,, 2025). Pannatier et al., (2024) describe a speculative sampling procedure but its validity is not established and the draft model is not described. An any-order causal transformer baseline reported by Hoogeboom et al., (2022) performed poorly in the MDMs context. We improve on the pure any-order causal model by letting it re-use computations from a powerful non-causal transformer.
Guo and Ermon, (2025) perform self-speculative decoding within an any-subset AR model, which, given an arbitrarily located prompt, in-fills missing tokens in a left-to-right fashion. In contrast, our model can condition on an arbitrarily located prompt and in-fill using any ordering too, more akin to standard MDMs. They rely on the XLNet architecture which faces optimization difficulties for generation tasks with short prompts and large numbers of missing tokens requiring optimization tuning such as a masking schedule warm-up (Yang et al.,, 2019). Our architecture does not require any mask schedule warm-up and can operate in the fully masked regime with no prompt.
In the MDMs context, previous works have also aimed at speeding up sampling. Similar to our work, some aim to introduce dependencies between the newly sampled tokens via an energy-based model (Xu et al.,, 2025) or a latent variable mixture model (Hayakawa et al.,, 2025). We instead opt for speculative sampling to sample from a non-factorized target in one shot. Liu et al., (2025) use an AR model to derive a target, however, their predictive distribution must be sampled autoregressively for each denoising step. Deschenaux and Gulcehre, (2025) use distillation to speed up sampling, however, their student model remains factorized, limiting potential speedups.
@@ 5.1 Text8
To measure sample quality we calculate spelling accuracy by generating $D=256$ characters and then computing the proportion of words within the sample that also appear in the training dataset; a word being defined as any number of lowercase characters between two whitespace characters. To create the accuracy-NFE tradeoff, we vary the window size and number of draft-verify steps (each requiring one forward pass of the causal block) undertaken per pass of the non-causal blocks. We utilize a cosine shaped window (similar to the cosine discretization grid proposed in Shi et al., 2024) with varying time step parameter, the derivation of which is given in Appendix D.
We compare against MDM using the implementation from Shi et al., (2024) with a varying number of timesteps used to simulate the generative process. We plot our results in Figure 3. We see that our approach is able to achieve higher spelling accuracy at lower NFE versus standard MDMs. For example in the low NFE range, our approach is able to achieve greater than $2\times$ reduction in NFE.
@@ 5.2 OpenWebText
We demonstrate our method on OpenWebText using a GPT2-scale, 150M parameter, 12-layer transformer with RoPE positional encodings (Shi et al.,, 2024), setting the first 11 layers as non-causal and the final layer as causal. We use generative perplexity (measured by GPT2) as our metric, supplementing it with unigram token entropy to ensure sample diversity (as generative perplexity can be cheated by low-temperature sampling). Compared to an MDM baseline (Shi et al.,, 2024) with an identical architecture, our method (Table 1) achieves the same generative perplexity with half the NFE, while maintaining similar sample entropy for all NFE levels.
We also compare with Self-Distillation Through Time (SDTT) (Deschenaux and Gulcehre,, 2025), where a student model is trained on a coarse time grid to match a teacher MDM sampled on a fine time grid. SDTT achieves very low GPT2 NLLs, far below the teacher MDM model using much larger NFE. Investigating the entropy of the produced samples reveals that SDTT samples are lower entropy than other methods. This mode-seeking is likely caused by truncation errors in the SDTT teacher model sampling (Zheng et al.,, 2025). In contrast, our approach reduces NFE while maintaining sample diversity equivalent to the baseline.
Table 1: GPT2 NLLs (in nats per dim) and unigram token entropy (nats). For each method a metric-NFE tradeoff curve is created by varying sampling parameters. Values at each NFE are read off by linearly interpolating between the two nearest points. For comparison we include results for mask diffusion and Self Distillation Through Time (SDTT) (Deschenaux and Gulcehre,, 2025). We also include ablations, removing the residual output connection and a 10 non-causal 2 causal block model.
@@ 5.3 Protein Sequence Modelling
We now demonstrate our method on non-text data by modeling amino acid sequences from the Uniref50 dataset, following settings from Wang et al., (2024) (40M sequences, max length 1022). To demonstrate applicability to pretrained models, we take the 150M parameter, 30-layer ESM2-based model from Wang et al., (2024) (Lin et al.,, 2023), freeze its parameters, and add a single causal block (carrying hidden states from the final non-causal block), training only this extra head.
We measure sample quality using the average pLDDT from ESMFold (Lin et al.,, 2023); a higher pLDDT indicates higher folding confidence, suggesting the sequence better follows the natural distribution. We compare this to the original Wang et al., (2024) non-causal model, sampled using the standard MDM algorithm (omitting resampling, temperature scaling, or confidence scoring for a consistent baseline). Our results (Figure 4) show a better sample quality-NFE trade-off than the standard algorithm, in particular achieving a $\sim 2\times$ speed-up for high pLDDT. This showcases the benefit of adding just a single causal block to a frozen pretrained network.
@@ 6 Discussion
We presented self-speculative masked diffusions, a novel architecture and sampling scheme for MDMs that reduces the number of network forward passes required for a given sample quality compared to standard MDM approaches. On a variety of datasets, from text to protein sequences, our method consistently improves computational efficiency, reducing the number of network evaluations by up to  $2\times$ compared to MDMs. Further work could explore the natural combination of our method with compute-intensive inference-scaling techniques, such as re-masking corrector steps, in order to reach a new level of model reasoning capability for a fixed compute budget.
@@
This appendix is organized as follows. In Section A, we present an illustration of the different attention mechanisms considered in this paper. Section B details the complete speculative sampling algorithm, including the windowing technique used in our experiments. Section C then provides proofs for all of our propositions, deriving the likelihood of our model when sampled using Algorithm 2 and deriving the distribution over the number of rejection steps within Algorithm 2. In Section D we discuss choices for the window function used during sampling including the derivation of a cosine style window. Finally, details for all of our experiments are provided in Section E.

### v2

@@ 3.3 Sampling Procedure
We sample our hybrid architecture using Algorithm 2. The general procedure follows Algorithm 1, revealing tokens in update steps until the sequence is complete. In a given update step (with $i$ tokens revealed), the non-causal blocks first generate draft tokens for all unknown positions, $\hat{\bm{x}}^{\sigma(i+1:D)}\sim\overset{\leftrightarrow}{p}_{\theta}(\hat{\bm{x}}^{\sigma(i+1:D)}|\theta(\bm{x}^{\sigma(1:i)}))$. The causal blocks then give target probabilities for the drafted tokens $\{\overset{\rightarrow}{p}_{\theta,\phi}(\hat{\bm{x}}^{\sigma(d)}|\theta(\bm{x}^{\sigma(1:i)}),\phi(\hat{\bm{x}}^{\sigma(i+1:d-1)}))\}_{d=i+1}^{D}$. Finally, a speculative sampling inner loop accepts some of the drafted tokens. Appendix B details a more general windowed procedure.
@@
Our procedure bears some similarity to self-speculative (Zhang et al., 2024; Elhoushi et al., 2024; Liu et al., 2024) and Medusa (Cai et al., 2024) style approaches where an initial subset of the network is used to parameterize a draft distribution which is then verified by the full transformer. However, using self-speculation for MDMs causes additional complexity because the target distribution given by Equation 6 changes depending on the number of revealed tokens. Consider the targets $\overset{\rightarrow}{p}_{\theta,\phi}(\bm{x}^{\sigma(i+1:D)}|\theta(\bm{x}^{\sigma(1:i)}),\phi)$ and $\overset{\rightarrow}{p}_{\theta,\phi}(\bm{x}^{\sigma(j+1:D)}|\theta(\bm{x}^{\sigma(1:j)}),\phi)$ for $j>i$. The factors for a fixed $d$ differ; i.e., importantly, we have $\overset{\rightarrow}{p}_{\theta,\phi}(\bm{x}^{\sigma(d)}|\theta(\bm{x}^{\sigma(1:i)}),\phi(\bm{x}^{\sigma(i+1:d-1)}))\neq\overset{\rightarrow}{p}_{\theta,\phi}(\bm{x}^{\sigma(d)}|\theta(\bm{x}^{\sigma(1:j)}),\phi(\bm{x}^{\sigma(j+1:d-1)}))$. This is because the non-causal blocks operate on different inputs (one sequence with $i$ revealed tokens, the other with $j$). So the hidden states passed to the causal blocks differ, causing the target distribution for position $\sigma(d)$ to change throughout the generation trajectory. This is not an issue for existing self-speculative models restricted to a single left-to-right ordering. All transformer blocks are causal in this case and never observe future tokens, meaning the target remains unchanged during generation.
@@ 3.4 Theoretical Results
Let $p_{\theta,\phi}(\bm{x}^{\sigma(1:D)}|\sigma)$ be the probability of generating $\bm{x}^{\sigma(1:D)}$ in the ordering $\sigma$ under Algorithm 2. It is non-trivial to reason about $p_{\theta,\phi}(\bm{x}^{\sigma(1:D)}|\sigma)$ because the inner loop target distribution changes whenever a rejection occurs. Furthermore, it is the alignment between $\overset{\leftrightarrow}{p}_{\theta}$ and $\overset{\rightarrow}{p}_{\theta,\phi}$ that determines the rejection probability and so both $\overset{\leftrightarrow}{p}_{\theta}$ and $\overset{\rightarrow}{p}_{\theta,\phi}$ influence $p_{\theta,\phi}(\bm{x}^{\sigma(1:D)}|\sigma)$. Therefore, to calculate $p_{\theta,\phi}(\bm{x}^{\sigma(1:D)}|\sigma)$, it would seem that one would have to sum the contributions of $\overset{\leftrightarrow}{p}_{\theta}$ and $\overset{\rightarrow}{p}_{\theta,\phi}$ over all combinations of possible numbers of tokens revealed during each inner loop which would be intractable for $D>>1$. Instead, we show in Proposition 3.1 that $p_{\theta,\phi}(\bm{x}^{\sigma(1:D)}|\sigma)$ can be tractably calculated with $D$ forward passes and $O(D^{2})$ operations using a recursive decomposition. The proof is provided in Appendix C.1.
@@
Consider the sampling scheme defined by Algorithm 2. For a given ordering $\sigma$, let $A^{\sigma(d)}$ denote the event that the token in position $\sigma(d)$ was accepted and let $R^{\sigma(d)}$ denote the event it was rejected and resampled. The distribution of samples output by Algorithm 2, $p_{\theta,\phi}(\bm{x}^{\sigma(1:D)}|\sigma)$, then admits the following decomposition (with $A^{\sigma(D+1:D)}:=\emptyset,~p_{\theta,\phi}(x^{\sigma(1:0)},R^{\sigma(0)}):=1$)
@@
Proposition 3.1 is a dynamic programming recursion where the total likelihood $p_{\theta,\phi}(\bm{x}^{\sigma(1:D)}|\sigma)$ is split up using the simpler object $p_{\theta,\phi}(\bm{x}^{\sigma(1:d)},R^{\sigma(d)})$ which can be written as a recursion on its previous values $p_{\theta,\phi}(\bm{x}^{\sigma(1:k-1)},R^{\sigma(k-1)})$ for $k=1$ to $d$. With a tractable expression for $p_{\theta,\phi}(\bm{x}^{\sigma(1:D)}|\sigma)$, we can now obtain an Evidence Lower BOund (ELBO) on $\log p_{\theta,\phi}(\bm{x})$,
@@ 3.5 Sampling Improvements
Algorithm 2 assumes each sampling inner loop corresponds to one forward pass of the non-causal blocks and a single verification pass through the remaining causal blocks. We can vary the sample quality-efficiency trade-off by running the speculative inner loop multiple times per single forward pass of the non-causal blocks. The non-causal forward pass first creates a draft $\overset{\leftrightarrow}{p}_{\theta}(\bm{x}^{\sigma(i+1:D)}|\theta(\bm{x}^{\sigma(1:i)}))$ distribution. Draft tokens are then sampled and we compute the target probabilities on these draft tokens using the causal layers. One speculative sampling loop is run to accept a number of these draft tokens finishing with a rejection and a resampling of a token. In the next speculative sampling inner loop we re-use the non-causal hidden states but recompute the target probabilities on the remaining draft tokens, noting the resampling operation will change the target probabilities for subsequent tokens compared to the first speculative sampling inner loop. This procedure is repeated for a user specified number of inner loops, the full algorithm is described in Appendix B. In our experiments, the vast majority of the network is non-causal, so this procedure greatly increases efficiency. To know when to stop generating tokens with the same draft $\overset{\leftrightarrow}{p}_{\theta}(\bm{x}^{\sigma(i+1:D)}|\theta(\bm{x}^{\sigma(1:i)}))$, we use windows (Leviathan et al., 2023; Chen et al., 2023) to set a maximum number of tokens we can reveal. We discuss window choices in Appendix D.
@@ 4 Related Work
Several speculative sampling works have explored using a single network for both the draft model and target model (Zhang et al., 2024; Elhoushi et al., 2024; Liu et al., 2024), termed self-speculative decoding. These differ from our work by using causal draft and target layers in a strict left-to-right setting. In the same setting, non-causal factorized predictions were explored by Cai et al. (2024) using a mostly causal stack with single-layer factorized prediction heads to draft tokens. We flip this: our network is mostly non-causal with a small final causal layer. Finally, a non-causal draft model and entirely separate causal target model was explored for audio generation in Ziv et al. (2024). We combine these two into a single architecture reducing deployment complexity.
Causal models for any-order AR training were explored by Yang et al. (2019), who introduced XLNet, a specialized two-stream architecture enforcing correct causal dependencies for different orderings. Further architectures used a standard single stream causal architecture and input ordering information through double positional encodings (Pannatier et al., 2024) or extra position tokens (Pang et al., 2025). Pannatier et al. (2024) describe a speculative sampling procedure but its validity is not established and the draft model is not described. An any-order causal transformer baseline reported by Hoogeboom et al. (2022) performed poorly in the MDMs context. We improve on the pure any-order causal model by letting it re-use computations from a powerful non-causal transformer.
Guo and Ermon (2025) perform self-speculative decoding within an any-subset AR model, which, given an arbitrarily located prompt, in-fills missing tokens in a left-to-right fashion. In contrast, our model can condition on an arbitrarily located prompt and in-fill using any ordering too, more akin to standard MDMs. They rely on the XLNet architecture which faces optimization difficulties for generation tasks with short prompts and large numbers of missing tokens requiring optimization tuning such as a masking schedule warm-up (Yang et al., 2019). Our architecture does not require any mask schedule warm-up and can operate in the fully masked regime with no prompt.
In the MDMs context, previous works have also aimed at speeding up sampling. Similar to our work, some aim to introduce dependencies between the newly sampled tokens via an energy-based model (Xu et al., 2025) or a latent variable mixture model (Hayakawa et al., 2025). We instead opt for speculative sampling to sample from a non-factorized target in one shot. Liu et al. (2025) use an AR model to derive a target, however, their predictive distribution must be sampled autoregressively for each denoising step. Deschenaux and Gulcehre (2025) use distillation to speed up sampling, however, their student model remains factorized, limiting potential speedups.
@@ 5.1 Text8
To measure sample quality we calculate spelling accuracy by generating $D=256$ characters and then computing the proportion of words within the sample that also appear in the training dataset; a word being defined as any number of lowercase characters between two whitespace characters. To create the accuracy-NFE tradeoff, we vary the window size and number of draft-verify steps (each requiring one forward pass of the causal block) undertaken per pass of the non-causal blocks. We utilize a cosine shaped window (similar to the cosine discretization grid proposed in Shi et al. 2024) with varying time step parameter, the derivation of which is given in Appendix D.
We compare against MDM using the implementation from Shi et al. (2024) with a varying number of timesteps used to simulate the generative process. We plot our results in Figure 3. We see that our approach is able to achieve higher spelling accuracy at lower NFE versus standard MDMs. For example in the low NFE range, our approach is able to achieve greater than $2\times$ reduction in NFE.
@@ 5.2 OpenWebText
We demonstrate our method on OpenWebText using a GPT2-scale, 150M parameter, 12-layer transformer with RoPE positional encodings (Shi et al., 2024), setting the first 11 layers as non-causal and the final layer as causal. We use generative perplexity (measured by GPT2) as our metric, supplementing it with unigram token entropy to ensure sample diversity (as generative perplexity can be cheated by low-temperature sampling). Compared to an MDM baseline (Shi et al., 2024) with an identical architecture, our method (Table 1) achieves the same generative perplexity with half the NFE, while maintaining similar sample entropy for all NFE levels.
We also compare with Self-Distillation Through Time (SDTT) (Deschenaux and Gulcehre, 2025), where a student model is trained on a coarse time grid to match a teacher MDM sampled on a fine time grid. SDTT achieves very low GPT2 NLLs, far below the teacher MDM model using much larger NFE. Investigating the entropy of the produced samples reveals that SDTT samples are lower entropy than other methods. This mode-seeking is likely caused by truncation errors in the SDTT teacher model sampling (Zheng et al., 2025). In contrast, our approach reduces NFE while maintaining sample diversity equivalent to the baseline.
Table 1: GPT2 NLLs (in nats per dim) and unigram token entropy (nats). For each method a metric-NFE tradeoff curve is created by varying sampling parameters. Values at each NFE are read off by linearly interpolating between the two nearest points. For comparison we include results for mask diffusion and Self Distillation Through Time (SDTT) (Deschenaux and Gulcehre, 2025). We also include ablations, removing the residual output connection and a 10 non-causal 2 causal block model.
@@ 5.3 Protein Sequence Modelling
We now demonstrate our method on non-text data by modeling amino acid sequences from the UniRef50 dataset, following settings from Wang et al. (2024) (40M sequences, max length 1022). To demonstrate applicability to pretrained models, we take the 150M parameter, 30-layer ESM2-based model from Wang et al. (2024) (Lin et al., 2023), freeze its parameters and add a single causal block (carrying hidden states from the final non-causal block), training only this extra head.
@@
We measure sample quality using the average pLDDT from ESMFold (Lin et al., 2023); a higher pLDDT indicates higher folding confidence, suggesting the sequence better follows the natural distribution. We compare this to the original Wang et al. (2024) non-causal model, sampled using the standard MDM algorithm (omitting resampling, temperature scaling, or confidence scoring for a consistent baseline). Our results (Figure 4) show a better sample quality-NFE trade-off than the standard algorithm, in particular achieving a $\sim 2\times$ speed-up for high pLDDT. This showcases the benefit of adding just a single causal block to a frozen pretrained network. We note that algorithms that also choose $\sigma$ can achieve higher pLDDT values, for example Peng et al. (2025), however we here focus on the standard MDM formulation with $\sigma$ selected uniformly at random.
@@ 6 Discussion
We presented self-speculative masked diffusions, a novel architecture and sampling scheme for MDMs that reduces the number of network forward passes required for a given sample quality compared to standard MDM approaches. On a variety of datasets, from text to protein sequences, our method consistently improves computational efficiency, reducing the number of network evaluations by up to  $2\times$ compared to MDMs while adding only a negligible computational overhead (a 0.98% increase in FLOPs in standard settings). Further work could explore the natural combination of our method with compute-intensive inference-scaling techniques, such as re-masking corrector steps, in order to reach a new level of model reasoning capability for a fixed compute budget.
@@
This appendix is organized as follows. In Section A, we present an illustration of the different attention mechanisms considered in this paper. Section B details the complete speculative sampling algorithm, including the windowing technique used in our experiments. Section C then provides proofs for all of our propositions, deriving the likelihood of our model when sampled using Algorithm 2 and deriving the distribution over the number of rejection steps within Algorithm 2. In Section D we discuss choices for the window function used during sampling including the derivation of a cosine style window. In Section E we provide a FLOP analysis of the additional overhead of our self-speculative architecture and in Section F we discuss the influence of hyperparameters on our method. Finally, details for all of our experiments are provided in Section G.